# Infinitely many genus two and three curves of small fixed positive rank

### v2

@@ 3. Genus three curves
We apply Theorem 3.1 for the following values of $a,b,c,d$. Here we list examples for $0\leq r\leq 4$ in the table provided below. All the algebraic ranks can be checked using Magma [BCP97].
$r$ $a$ $b$ $c$ $d$
0 1 2 3 8
1 1 2 3 9
2 1 2 3 36
3 -50 -5 42 50
4 -50 1 27 50
4. Rational points on such curves and a challenge
4.1. Rational points on these curves
In this section, we only consider curves defined over $\mathbb{Q}$, so we exclude those from §2.4. Even though we construct curves of higher rank, it is easy to determine their rational points in most of our examples (except for $C_{p,d}$ from §2.2 when $p\equiv 3\pmod{4}$) because these curves admit a quotient map to a rank 0 elliptic curve. This elliptic curve has only finitely many points, and we can determine which pullbacks of these points are rational points. Now we focus on all but finitely many curves, which could still have more rational points.
On the other hand, also by [Sil09, Exercise 10.19], for $E_{p}\colon y^{2}=x^{3}+p^{3}$ with $p\equiv 5\pmod{12}$, we have $E_{p}(\mathbb{Q})=\{O,(-p,0)\}$. Hence, if $d$ in a square (i.e. $d=1$ since we assume $d$ is squarefree), then $C_{p,d}(\mathbb{Q})=\{\infty_{\pm}\}$. Otherwise, the set $C_{p,d}(\mathbb{Q})$ is empty. When $p\equiv 3\pmod{4}$, the problem of determining $C_{p,d}(\mathbb{Q})$ could be very difficult.
In Section 3, the rank 0 family of twists of elliptic curves is $E_{D}$ (using the notation from Section 3), for which we know $E_{D}(\mathbb{Q})[2]\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ and that it is a change of variables of $E^{\prime}_{D}$. Since twisting will not preserve rational coordinates of an $n$-torsion point, unless $n=2$, for almost all $D$, we will have that $E_{D}(\mathbb{Q})=E_{D}(\mathbb{Q})[2]$ (again, we recall that we consider such curves of rank 0). Pullback of these points correspond to the roots of the defining polynomial in $x$ of $E^{\prime}_{D}$ as well as those of $C_{3,D}$. Since we vary $D$, for almost all $D$ none of the $aD$, $bD$, $cD$, and $dD$ will be a square for fixed $a,b,c,d\in\mathbb{Z}$. We have $C_{3,D}(\mathbb{Q})=\emptyset$ almost always.
4.2. Challenge
@@
If the Jacobian of a curve is decomposable as a product of Jacobians of curves, then we reduce the problem of computing the rank to smaller genus curves, which is, in principle, way easier. We believe that the problem becomes much harder if we restrict the question for curves $C$ of genus two such that $\mathrm{Jac}(C)$ is absolutely simple, which gives the final question.
@@ Acknowledgements
We would like to thank the Max Planck Institute for Mathematics for its generous support. We would like to thank Jennifer Balakrishnan, Dušan Dragutinović, Peter Koymans, Oana Padurariu, Carlo Pagano, Lazar Radičević, Alex Smith, and Wadim Zudilin for helpful discussions.

### v3

@@ 3.1. Higher genus - over any number field
We apply Theorem 3.8 to any $B$-generic hyperelliptic curve $C$ of genus $g$, which can easily be constructed using the Chinese Remainder Theorem. ∎
Theorem 3.10.
Let $g\geq 3$ be a positive integer and denote $k=\lfloor\frac{g+1}{2}\rfloor$. Assume there are $b_{1},\ldots,b_{2k+1}\in\mathcal{O}_{K}$ so that the hyperelliptic curve
$X\colon y^{2}=(x-b_{1})\cdots(x-b_{2k+1})$
has $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$. Suppose for each specified value of $g$ as mentioned below, the curve $Y_{1}$ satisfies Condition 3.5:
  • $g=2k$, and $Y_{1}:y^{2}=x(x-b_{1})\cdots(x-b_{2k+1})$.
  • g=2k1g=2k-1, and $Y_{1}:y^{2}=(x-(b_{1}-b_{2k+1}))\cdots(x-(b_{2k}-b_{2k+1}))$.
Then, there are infinitely many hyperelliptic curves of genus $g$ whose Jacobian over $K$ has rank $r$.
Proof.
Suppose $g=2k$, Let $Y_{d}\colon dy^{2}=x(x-b_{1})\cdots(x-b_{2k+1})$ be a quadratic twist of $Y_{1}$ by $d$. By Theorem 3.8, there are infinitely many such quadratic twists $Y_{d}$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$. Hence, there are infinitely many genus $g$ curves
$C_{d}\colon dy^{2}=(x^{2}-db_{1})\cdots(x^{2}-db_{2k+1})$
of rank $r$ as $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ by Lemma 3.1.
Similarly, if g=2k1g=2k-1, then,
$X\cong X^{\prime}\colon y^{2}=x(x-(b_{1}-b_{2k+1}))\cdots(x-(b_{2k}-b_{2k+1})).$
By Theorem 3.8, there are infinitely many twists
$Y_{d}\colon dy^{2}=(x-(b_{1}-b_{2k+1}))\cdots(x-(b_{2k}-b_{2k+1}))$
of rank 0. By Lemma 3.1, there are infinitely many curves
$C_{d}\colon dy^{2}=(x^{2}-d(b_{1}-b_{2k+1}))\cdots(x^{2}-d(b_{2k}-b_{2k+1}))$
of genus $g$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=r$. ∎
Remark 3.11.
We note that the affine models for curves $Y_{1}$ in both cases $g=2k$ and g=2k1g=2k-1 are defined in terms of even degree polynomials in $x$. For computational purposes, we rewrite the models for $Y_{1}$ in terms of odd degree polynomials in $x$. For the case when $g=2k$, let $\mathcal{B}:=\prod_{i=1}^{2k+1}b_{i}$. Then the equation defining $Y_{1}$ can be rewritten as
$Y_{1}\cong Y_{1}^{\prime}:y^{2}=\left(x+\frac{\mathcal{B}}{b_{1}}\right)\cdot\left(x+\frac{\mathcal{B}}{b_{2}}\right)\cdots\left(x+\frac{\mathcal{B}}{b_{2k+1}}\right).$
If g=2k1g=2k-1, let $\mathcal{B}:=\prod_{i=2}^{2k}(b_{i}-b_{1})$. Then we can rewrite the equation defining $Y_{1}$ as
$Y_{1}\cong Y_{1}^{\prime}:y^{2}=\left(x+\frac{\mathcal{B}^{\prime}}{b_{2}-b_{1}}\right)\cdot\left(x+\frac{\mathcal{B}^{\prime}}{b_{3}-b_{1}}\right)\cdots\left(x+\frac{\mathcal{B}^{\prime}}{b_{2k}-b_{1}}\right).$
Then, in our applications, see §3.2, we check Condition 3.5 directly for these odd degree models using Magma.
We can construct such examples in a different way, i.e., coming of the hyperelliptic curves, whose Jacobians have very few $K$-rational 2-torsion points, using the result of Yu, [Yu19, Theorem 1.2].
Theorem 3.12.
Let $g\geq 3,r\geq 0$ be integers and $K$ a number field.
  1. (1)
    Assume that $g=2k$ for some $k\in\mathbb{N}$ and that there exists a hyperelliptic curve $X:y^{2}=f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, where $f\in\mathcal{O}_{K}[x]$, $\deg(f)=2k+1$, is a monic polynomial whose splitting field has Galois group isomorphic to $S_{2k+1}$. Suppose in addition that there exists a square-free element $m\in\mathcal{O}_{K}$ such that the curve $Y_{m}:my^{2}=xf(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$. Then, there are infinitely many genus $g$ curves $C_{d}:dy^{2}=f_{d}(x^{2})$ of rank $r$, where $d\in\mathcal{O}_{K}$.
  2. (2)
    Assume that g=2k1g=2k-1 for some $k\in\mathbb{N}$ and that there exists a hyperelliptic curve $X:y^{2}=x(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, where $f\in\mathcal{O}_{K}[x]$, $\deg(f)=2k-1$, is a monic polynomial whose splitting field has Galois group isomorphic to $S_{2k-1}$. Suppose in addition that there exists a square-free element $m\in\mathcal{O}_{K}$ such that the curve $Y_{m}:my^{2}=(x-1)f(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$. Then, there are infinitely many many genus $g$ curves $C_{d}:dy^{2}=(x^{2}-d^{2})f_{d}(x^{2})$ of rank $r$, where $d\in\mathcal{O}_{K}$.
Proof.
  1. (1)
    We note that the existence of a square-free element $m\in\mathcal{O}_{K}$ such that $Y_{m}:my^{2}=xf(x)$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(Y_{m})/K)\equiv 0\mod 2$ implies that the disparity $\delta(\mathrm{Jac}(Y_{1})/K)\neq-\frac{1}{2}$ appearing in the statement of [Yu19, Theorem 1.2]. The assumptions assure that we can apply [Yu19, Theorem 1.2] to ensure that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=xf(x)$ of $Y\colon y^{2}=xf(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$. Then, the statement follows from Lemma 3.1 which gives that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$, and the argument that there are infinitely many curves that are non-isomorphic among them is as before.
  2. (2)
    The same as (1), and now we apply [Yu19, Theorem 1.2] to conclude that there are infinitely many quadratic twists $Y_{d}\colon dy^{2}=(x-1)f(x)$ of $Y\colon y^{2}=(x-1)f(x)$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=0$.
Theorem 3.13.
Let $k$ be a positive integer. There is $B>0$ such that the following holds. Let $X:y^{2}=(x-a_{1})\cdots(x-a_{2k+1})$ be a hyperelliptic curve over a number field $K$ that is BB-über-generic such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$. Then there are infinitely many hyperelliptic curves $X/K$ of genus $g$, where $g\in\{2k-1,2k\}$ satisfying $\mathrm{Rank}(\mathrm{Jac}(C)(X))=r+1$.
Proof.
The easier case is when g=2k1g=2k-1. We may assume that $a_{2k+1}=0$ (otherwise just replace $x-a_{2k+1}$ by $x$). We need to glue the curve $C$ with $Y:y^{2}=(x-a_{1})\cdots(x-a_{2k})$. We choose $B>0$ from [KM, Theorem 1.3] for genus k1k-1. Our assumption of being BB-über-generic clearly implies that $Y$ is $B$-generic. Hence, by [KM, Theorem 1.3], there are infinitely many twists $Y_{d}:dy^{2}=(x-a_{1})\cdots(x-a_{2k})$ of $Y$ such that $\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))=1$. Then, for $C_{d}:dy^{2}=(x^{2}-da_{1})\cdots(x^{2}-da_{2k})$, by Lemma 3.1, we have that $\mathrm{Jac}(C_{d})$ is isogenous to $\mathrm{Jac}(X)\times\mathrm{Jac}(Y_{d})$ over $K$. Hence, $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=r+1$, and the argument that there are infinitely many non-isomorphic curves among them is as before.
Now, we consider the case $g=2k$. We may assume now $a_{1}\cdots a_{2k+1}\neq 0$. We need to glue the curve $C$ with $Y:y^{2}=x(x-a_{1})\cdots(x-a_{2k+1})$. We choose $B>0$ from [KM, Theorem 1.3] for genus $k$. We need an odd model of $Y$, so the change of variables $x\mapsto\frac{1}{x}$, induces that $Y$ is isomorphic over $K$ to
$Z:y^{2}=\left(x+\frac{A}{a_{1}}\right)\cdots\left(x+\frac{A}{a_{2k+1}}\right),$
where $A=a_{1}\cdots a_{2k+1}$. Now, the assumption that $X$ is BB-über-generic gives that for any $i,j$ and valuation $w$
$\mathrm{val}_{w}\left(\frac{A}{a_{i}}-\frac{A}{a_{j}}\right)=\mathrm{val}_{w}\left(\frac{A}{a_{i}a_{j}}(a_{j}-a_{i})\right)=\mathrm{val}_{w}(a_{j}-a_{i}),$
so $Z$ is $B$-generic. Therefore, we can apply [KM, Theorem 1.3] to conclude that there are infinitely many twists $Z_{d}$ of $Z$ such that $\mathrm{Rank}(\mathrm{Jac}(Z_{d})(K))=1=\mathrm{Rank}(\mathrm{Jac}(Y_{d})(K))$ for the corresponding twist $Y_{d}$ of $Y$. Now, we conclude as in the previous case. ∎
Remark 3.14.
Using the same idea of the proof that $X$ being BB-über-generic implies $Y_{d}$ being $B$-generic, we can use Remark 3.6 and [Smi22b, Theorem 2.14] to prove that if $X$ is BB-über-generic and $\mathrm{Rank}(\mathrm{Jac}(X)(K))=r$, then there are infinitely many hyperelliptic curves of genus $g$ whose Jacobian over $K$ has rank $r$.
Theorem 3.15.
Let $g\geq 3$ be an integer and $K$ a number field. There are infinitely many hyperelliptic curves $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=r$ for $r=1$ and $2$.
Proof.
By [KM], we can always find a quadratic twist of a BB-über-generic curve $X$ of genus $\lfloor\frac{g+1}{2}\rfloor$ whose Jacobian over $K$ has rank $1$. We then apply Theorem 3.10 (++ Remark 3.6) and Theorem 3.13 to conclude. ∎
3.2. Examples over $\mathbb{Q}$
Now we prove the existence of infinitely many curves over $\mathbb{Q}$ of genus $g$ whose Jacobian over $\mathbb{Q}$ as rank $r$ for $g\in\{3,4\}$ and $1\leq r\leq 4$ and $g\in\{5,6\}$ and $1\leq r\leq 3$. Even though the existence of the rank 1 and 2 follows from Theorem 3.15, we still decided to include these examples.
We will always use Theorem 3.10. We will find a hyperelliptic curve of the shape
$X\colon y^{2}=(x-b_{1})\cdots(x-b_{2k+1})$
such that $\mathrm{Rank}(\mathrm{Jac}(X)(\mathbb{Q}))=r$ and check that the conditions for the curve $Y_{1}$ from Theorem 3.10 using Remark 3.7. We note that only in the case for genus three curves, we do not need to do any checks because in this case $Y_{1}$ is an elliptic curve and then the result about twisting comes automatically from [Smi25].
3.2.1. Genus three and four
We present a table of genus two curves $X\colon y^{2}=(x-b_{1})\cdots(x-b_{5})$, which, by Remark 3.7 and Theorem 3.10, prove the existence of infinitely many hyperelliptic curves $C/\mathbb{Q}$ of genus three and four such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$, for the following $r$ from the table.
$r$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$ $b_{5}$
1 1 2 3 4 10
2 1 2 3 4 37
3 -49 -4 1 43 51
4 -49 1 2 28 51
3.2.2. Genus five and six
We present a table of genus three curves $X\colon y^{2}=(x-b_{1})\cdots(x-b_{7})$, which, by Remark 3.7 and Theorem 3.10, prove the existence of infinitely many hyperelliptic curves $C/\mathbb{Q}$ of genus five and six such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$, for the following $r$ from the table.
$r$ $b_{1}$ $b_{2}$ $b_{3}$ $b_{4}$ $b_{5}$ $b_{6}$ $b_{7}$
1 1 2 3 4 5 6 7
2 1 2 3 4 6 7 11
3 1 2 3 4 5 14 21
Remark 3.16.
Even though it sounds appealing to use Theorem 3.13 to construct examples, we note that the $B$-generic and BB-über-generic conditions are less practical to use. Namely, a careful inspection on the condition on $B$ from [KM] shows that $B>9600$ for genus two curves, so we will need to use several large primes to construct such $B$-generic or BB-übergeneric curves. Moreover, solving the Chinese Reminder Theorem system will result in large coefficients of the curve $C$. Hence, we have not tried to use $B$-generic or BB-übergeneric curves to get concrete examples.
Also, Theorem 3.12 is more tricky for applications because it is much easier to do explicit descent for elliptic curves defined by a polynomial which splits completely, rather than of those whose Galois group is as maximal as possible. Hence, we have not tried to use this theorem to construct examples either.
4. An answer to our question
@@
If the Jacobian of a curve is decomposable as a product of Jacobians of curves, then we reduce the problem of computing the rank to smaller genus curves, which is, in principle, way easier. The problem becomes much harder if we restrict the question for curves $C$ of any genus such that $\mathrm{Jac}(C)$ is absolutely simple. The groundbreaking work of Peter Koymans and Adam Morgan [KM] confirms that there are infinitely many absolutely simple Jacobians of hyperelliptic curves such that the rank of $\mathrm{Jac}(C)(\mathbb{Q})$ is equal to $1$. This result resolves the following question, which was posed as a challenge in the previous version of this article as stated in June 2025.
@@
We keep the question as stated in its original form, as [KM, Theorem 1.1] specifically cites the problem mentioned beforehand. We sincerely congratulate and thank them for solving the question stated above.
@@ Acknowledgements
We would like to thank the Max Planck Institute for Mathematics for its generous support. We would like to also thank Lodha Mathematical Sciences Institute for its generous support and hospitality during the Thematic Programme on Arithmetic Statistics where substantial results of this paper were achieved. We would especially like to thank Jennifer Balakrishnan, Maarten Derickx, Dušan Dragutinović, Peter Koymans, Adam Morgan, Carlo Pagano, and Lazar Radičević for numerous conversations and their enlightening comments about this work. We would like also to thank Jordan Ellenberg, Matija Kazalicki, Jef Laga, Filip Najman, Oana Padurariu, Alex Smith, and Wadim Zudilin for helpful discussions.